OpenStudy (anonymous):

Find the first, fourth, and tenth terms of the arithmetic sequence described by the given rule. a(n)=1+(n-1)(-5.7)

OpenStudy (anonymous):

anybody

OpenStudy (anonymous):

Shouldn't this be in the math section

OpenStudy (anonymous):

Sequences are like normal functions (the only difference is that the domain, in case you already know what it is, are the positive integers). So when you ask for the 1st, 4th and 10th elements, you only need to substitute these numbers into the equations, like this A(1)= -3 + (n-1)(-2.2) = -3 + (1-1)(-2.2)= -3 + 0 = -1 A(4)= -3 + (n-1)(-2.2) = -3 + (4-1)(-2.2) = A(10)= -3 + (n-1)(-2.2) = -3 + (10-1)(-2.2) =

OpenStudy (anonymous):

So, what do you think the next answer would be?

OpenStudy (anonymous):

O

OpenStudy (anonymous):

What do you think the next two answers are?